# MotionLab: Unified Human Motion Generation and Editing via the Motion-Condition-Motion Paradigm

### v3

@@ 5.2. Task Instruction Modulation
MM-DiT incorporates a modulation mechanism to enhance text-to-image generation by inputting text embedding as modulation. However, in our unified framework, tasks involve diverse and multiple modalities, making simple input strategies like task-specific modalities or a learned task token (e.g., [TASK]) insufficient to handle this complexity. Noticing the flexibility of natural language, we leverage text representations learned by foundation models (e.g., CLIP (Radford et al., 2021)) to enable task differentiation in MotionLab. For example, instead of directly inputting the corresponding style for motion style transfer, we use the text embedding of “edit source motion by given style”. This simple yet effective modulation approach enhances flexibility and scalability, allowing seamless extension to various tasks involving multiple modalities.
@@ 5.3. Motion Curriculum Learning
Pre-training. Intuitively, the reconstruction of masked source motion is the easiest task. Hence, we first train the model based on the masked source motion, independent of the conditions. This approach allows the model to learn prior motion representations independent of conditions, thereby generalizing to different tasks. For the masked ratio of source motion, we propose randomly masking from zero frames to all frames. Compared to the fixed 75% masked ratio like MAE (He et al., 2022), this flexible strategy provides tasks of varying difficulty levels, avoiding overfitting on simple tasks (all frames) and mode collapse (zero frames) on difficult tasks. Furthermore, this strategy seamlessly performs source motion reconstruction (i.e., all frames) and unconditional training (i.e., zero frames), which is crucial for CFG in Section 5.4.
Additionally, motion in-between and trajectory-based generation can be viewed as generalized cases of masked reconstruction by masking certain frames or joints’ coordinates. Since they share the same condition path as discussed in Sec. 4, we remove the text condition and treat their source and target motions as additional reconstruction motion pairs for pre-training. Specifically, we pre-train MotionLab using these three sources for 1,000 epochs.
@@
Fine-tuning. In the supervised fine-tuning stage, we train MotionLab on tasks in an easy-to-hard sequence. Specifically, a new task is introduced into training every 200 epochs in the following order: ➀ text-based generation, ➁ style-based generation (an auxiliary task for training the condition path of the style, not our primary goal), ➂ trajectory-based editing (without text), ➃ text-based editing, ➄ style transfer, ➅ motion in-between and trajectory-based generation, ➆ trajectory-based editing. This progressive learning strategy ensures effective adaptation and knowledge sharing across tasks. Particularly, ➀ and ➁ are the simplest task because they only include one modality, whereas others include at least two modalities. Among tasks involving two modalities, ➂, ➃, and ➄ take priority over ➅ since they are editing tasks. Additionally, as text is less specific than trajectory but more specific than style, the order is ➂, ➃, and ➄.
@@
To mitigate catastrophic forgetting, previous tasks are trained alongside new tasks, based on the probability derived from the FID of the last evaluation. However, the FID scales for different tasks vary due to their differing difficulty levels. Consequently, we use the percentage change compared to the previous evaluation as the probability, which encourages the model to re-learn forgotten tasks or tasks that it has not yet fully mastered. To support classifier-free guidance, we also train the model to unconditionally generate and reconstruct the complete source motion. Empirically, in this stage, a 5% probability is allocated for unconditional generation, 5% for reconstructing the complete source motion, 45% for previous tasks, and 45% for the new task.
In summary, this training strategy, transitioning from a single modality to multiple modalities, enables the model to progressively learn each modality’s representation independently before capturing interactions between multiple modalities. This approach offers three key benefits: 1) adaptability to various tasks; 2) seamless support for CFG during inference; and 3) flexible training management, minimizing the need for retraining due to errors.
@@ 5.4. MotionLab Inference
Ours 0.223 0.805 9.589 2.881 3.029 0.068
Table 3. Evaluation of text-based motion generation on HumanML3D(Guo et al., 2022a) dataset. All AITS have been recalculated on RTX 4090D. The models in bold are the optimal models, and the models in underline are the sub-optimal models.
@@
Table 4. Evaluation of trajectory-based motion generation on HumanML3D(Guo et al., 2022a) dataset. AITS of all models have been recalculated on RTX 4090D.
@@
Table 5. Evaluation of text-based and trajectory-based motion editing on MotionFix (Athanasiou et al., 2024) dataset. TMED mean that we re-implement the models since original models are trained on the skeleton of SMPL format, while our models are trained on HumanML3D format. AITS of all models have been recalculated on RTX 4090D.
@@ 6. Experiments
Datasets. We evaluate our framework using the following datasets. To evaluate the text-based motion generation, the trajectory-based motion generation, motion in-between and motion style transfer, we leverage the HumanML3D (Guo et al., 2022a) dataset, which comprises 14646 motions and 44970 motion annotations. To evaluate on the text-based and trajectory-based motion editing, we utilize MotionFix (Athanasiou et al., 2024) dataset, which is the first dataset for text-based human motion editing including 6730 motion pairs.
@@
Evaluation Metrics. We evaluate our framework using the following metrics: 1) To evaluate the text-based motion generation, following the Chen et al., we introduce the FID to evaluate the distribution gap between the generated and original motions; Diversity to calculate the corresponding variance between motions; R-precision (R@K) to measure the proximity of the generated motion to the text or motion; Foot skating ratio to evaluate the physical plausibility of motion; Multi-modal Distance (MM Dist) calculates the distance between motions and texts. We also introduce we Average Inference Time per Sample (AITS) measured in seconds to evaluate the inference efficiency; 2) To evaluate the trajectory-based motion generation and motion in-between, following Xie et al., we introduce the Average Error to measures the mean distance between the generated motion locations and the keyframe locations; 3) To evaluate the text-based and trajectory-based motion editing, following Athanasiou et al., we introduce the AvgR to measure the success rate of retrieval from edited motion to target motion; 4) To evaluate the motion style transfer, following the Song et al., we introduce the Style Recognition Accuracy (SRA) and Content Recognition Accuracy (CRA) to measure the stylistic and content accuracy of the generated motion; Trajectory Similarity Index (TSI) to evaluate the trajectory preservation from source motion.
Method text gen. (FID) traj. gen. (avg. err.) text edit (R@1) traj. edit (R@1) in-between (avg. err.) style transfer (SRA)
w/o rectified flows 0.334 0.0359 54.38 69.21 0.0289 63.96
w/o MotionFlow Transformer 0.534 0.0447 51.26 65.34 0.0349 53.83
w/o Aligned ROPE 0.246 0.0886 45.39 61.99 0.0756 56.59
w/o task instruction modulation 0.279 0.0401 55.96 70.01 0.0288 63.91
w/o motion curriculum learning 2.236 0.1983 28.56 36.61 0.1682 34.23
Ours 0.209 0.0398 41.44 59.86 0.0371 67.55
Ours 0.223 0.0334 56.34 72.65 0.0273 64.97
Table 6. Ablation studies of our MotionLab’s main designs on each task. Ours means that we use this framework to train models in the same size for each task separately. For text-based motion generation, we compare FID; for trajectory-based motion generation and motion in-between. we compare average error; for motion editing, we compare R@1; for motion style transfer, we compare the SRA. Additional ablation studies are available in the supplementary.
@@
Implementation Details. In order to fairly compare our model with other models, motions from all datasets have been retargeted into one skeleton following HumanML3D format with 20 fps, where the number of joint $J$ is 22 and the dimension of motion feature $D$ is 263. The learning rate is set to be 1$\times$10-4. The timesteps are set to 1000 for training and 50 for inference. Our models are trained by four RTX 4090D with each batch of 64. To ensure a fair comparison, the AITS of all models are recalculated using one RTX 4090D.
@@ 6.1. Quantitative Results
Overall Performance. As shown in Table 3 to Table 5, and Figure 7 to Figure 8, MotionLab demonstrates promising performance across all benchmarks, underscoring the effectiveness of our framework’s design. Notably, as MotionLab is a unified framework without task-specific designs, it must balance versatility, performance, and efficiency. While it does not surpass the state-of-the-art in text-based motion generation (Table 3), it achieves competitive performance in this task and excels in other generation and editing tasks.
Specifically, as shown in Table 3 and Figure 10, MotionLab achieves superior performance (lowest FID, which is the key metric for generation tasks) with relatively fast inference time (third-lowest AITS). For trajectory-based tasks (Table 4) and the motion in-between task (Figure 7), MotionLab achieves lower average error. We believe these improvements stem from the effectiveness of masked pre-training and Aligned ROPE, which ensures spatial and temporal synchronization between the trajectory and target motion.
@@ 6.3. Ablation Studies
In this section, we perform several ablation experiments on our framework to validate the designs in MotionLab: the first variant replaces rectified flows with diffusion models; the second variant uses a regular transformer (i.e., without modulation mechanism and adopting cross-attention) instead of MFT. The third variant uses the implicit 1D-learnable encoding used in the past instead of Aligned ROPE; The fourth variant does not adopt the Task Instruction Modulation; the fifth variant directly learns all tasks based on their FID compared to the last evaluation. Meanwhile, we also use the same model to train specialist models for each task.
As illustrated in Table 6, the removal of motion curriculum learning markedly diminishes model performance across all tasks, underscoring its pivotal role in facilitating knowledge transfer between diverse tasks. This phenomenon may be attributed to the strategy’s capacity to enable the model to integrate its comprehension of spatial conditions (e.g., source motion, trajectory, and intermediate states) with abstract conditions (e.g., text and style), given that the latter can be partially represented by the former.
Also, as shown in Figure 9 and Table 5, Aligned ROPE is essential for time-related tasks, significantly reducing the average error. It effectively aligns source motion and target motion temporally, contributing to high R-precision in editing tasks. According to the results of Our, our unified framework achieves comparable outcomes to specialist models with the same parameters in each conditional path. Notably, our unified framework outperforms specialist models with a lower average error, likely due to the masked pre-training that aligns spatial information. Furthermore, our unified framework also surpasses specialist models in editing tasks, potentially due to the increased data from generation tasks and the learned conditional differences between source motion and target motion. Additionally, we validate our framework on the timesteps of inference. As shown in Figure 10, our framework demonstrates an optimal balance between generation quality and efficiency.
@@ 7. Conclusion
However, our work has several limitations for future improvement. Specifically, our framework primarily focuses on limb motion and does not account for finger or facial features. This limitation poses challenges for practical applications and necessitates additional processing by users. Future research may explore the integration of motion tasks involving fingers and faces into MotionLab.
@@ Appendix A Additional Quantitative Results
Figure 7. Comparison of the motion in-between with CondMDI (Cohan et al., 2024) on HumanML3D (Guo et al., 2022a), which shows that our model outperforms CondMDI.
@@ Appendix B Additional Ablation Studies
To further validate the Aligned ROPE, we also introduce the variant of 3D-Learnable and 3D-ROPE to distinguish the source motion, target motion and trajectory. As shown in Table 7, 1D-position encoding are better than 3D-position encoding by avoiding introducing distances between different modalities, and ROPE are better than learnable position encoding by explicit positional encoding. Hence our 1D-ROPE outperforms all others variants, demonstrating its effective to embed the position information into tokens.
To further validate the motion curriculum learning, we adopt the variant of removing the masked pre-training and directly supervised fine-tuning in order; the variant of with masked pre-training but supervised fine-tuning all tasks together; the variant of introducing masked reconstruction, motion in-between and trajectory based motion generation in orderly.
As shown in Table 8, our proposed motion curriculum learning outperforms all other variants, highlighting the effective of masked pre-training and fine-tuning tasks in order by avoiding gradient conflicts between different tasks. Specifically, the variant of masked pre-training in order demonstrates that necessity of introduce motion in-between and trajectory-based motion generation together, or will greatly weakens the performance of the model in the latter task.
Figure 9. Ablation results of MotionLab on the motion in-between. Beige motion is use 1D-learnable position encoding, purple motion use Aligned ROPE, and gray motions are the poses provided in keyframes, demonstrating the importance of Aligned ROPE.
Figure 10. Comparison of the inference time on text-based motion generation. We calculate AITS on the test set of HumanML3D (Guo et al., 2022a) without model or data loading parts. All tests are performed on the same RTX 4090D. The closer the model’s points are to the lower left corner, the stronger the model is.
Method text gen. (FID) traj. gen. (avg. err.) text edit (R@1) traj. edit (R@1) in-between (avg. err.) style transfer (SRA)
1D-Learnable 0.246 0.0886 45.39 61.99 0.0756 56.59
3D-Learnable 0.346 0.1865 35.46 53.74 0.1460 58.81
3D-ROPE 0.241 0.0579 51.34 70.00 0.0354 62.46
1D-ROPE (ours) 0.223 0.0334 56.34 72.65 0.0273 64.97
@@
Method text gen. (FID) traj. gen. (avg. err.) text edit (R@1) traj. edit (R@1) in-between (avg. err.) style transfer (SRA)
random selection based on FID 2.236 0.1983 28.56 36.61 0.1682 34.23
removing the masked pre-training 0.861 0.0932 44.99 63.92 0.0639 57.59
supervised fine-tuning all tasks together 1.331 0.1317 38.19 55.22 0.1143 50.59
masked pre-training in order 0.256 0.0423 56.33 69.31 0.0264 64.39
motion curriculum learning 0.223 0.0334 56.34 72.65 0.0273 64.97

### v4

@@ 5.2. Task Instruction Modulation
MM-DiT implements a modulation mechanism that enhances text-to-image generation through the incorporation of textual embeddings (e.g., “a photo of dog”) as modulation signals. However, within our unified framework, various tasks necessitate the integration of multiple modalities, and critically, identical modalities may require distinct representational forms across different tasks. This complexity renders approaches such as learned task tokens (e.g., [TASK]) or one-hot encoding vectors inadequate for managing arbitrary numbers and combinations of modalities.
Recognizing the inherent flexibility of natural language, we leverage textual representations acquired by foundation models (e.g., CLIP) to effectively differentiate identical modalities across disparate tasks. For instance, we utilize the textual embedding of “edit source motion by given style” to facilitate the adaptation of source motion to style transfer operations. This approach, while conceptually straightforward, provides remarkable effectiveness in enhancing system flexibility and scalability, thereby enabling seamless extension to diverse tasks involving multiple modalities.
@@ 5.3. Motion Curriculum Learning
Pre-training. Intuitively, the reconstruction of masked source motion is the easiest task. Hence, we first train the model based on the masked source motion, independent of the conditions. This approach allows the model to learn prior motion representations independent of conditions, thereby generalizing to different tasks. Following MoMask (Guo et al., 2024a), we randomly masking from zero frames to all frames. This flexible strategy provides tasks of varying difficulty levels, avoiding overfitting on simple tasks (all frames) and mode collapse on difficult tasks (zero frames). Furthermore, this strategy seamlessly performs source motion reconstruction (i.e., all frames) and unconditional training (i.e., zero frames), which is crucial for CFG. However, MoMask have to mask the coordinates and velocities of all joints in one frame simultaneously due to its discrete tokens. Therefore, we here extend the masked pre-training to randomly mask joints’ trajectories (not joints’ velocities for the user usage) for boosting the understanding of in-between and trajectory-based tasks. Specifically, we pre-train MotionLab using these three tasks (i.e., masked source motion reconstruction, trajectory-based generation without text and in-between without text) for 1,000 epochs.
@@
Fine-tuning. In the supervised fine-tuning stage, we train MotionLab on tasks in an easy-to-hard sequence. Specifically, a new task is introduced into training every 200 epochs in the following order: ➀ text-based generation, ➁ style-based generation (an auxiliary task for training the modality path of the style, not our primary goal), ➂ trajectory-based editing (without text), ➃ text-based editing, ➄ style transfer, ➅ motion in-between and trajectory-based generation, ➆ trajectory-based editing. This progressive learning strategy ensures effective adaptation and knowledge sharing across tasks. Particularly, ➀ and ➁ are the simplest task because they only include one modality, whereas others include at least two modalities. Among tasks involving two modalities, ➂, ➃, and ➄ take priority over ➅ since they are editing tasks. Additionally, as text is less specific than trajectory but more specific than style, the order is ➂, ➃, and ➄.
@@
To mitigate catastrophic forgetting, previous tasks are trained with new tasks, based on the probability derived from the FID of the last evaluation. However, the FID scales for different tasks vary due to their differing difficulty levels. Consequently, we use the percentage change compared to the previous evaluation as the probability, which encourages the model to re-learn forgotten tasks or tasks that it has not yet fully mastered. To support classifier-free guidance, we also train the model to unconditionally generate and reconstruct the complete source motion. Empirically, in this stage, a 5% probability is allocated for unconditional generation, 5% for reconstructing the complete source motion, 45% for previous tasks, and 45% for the new task.
In summary, this training strategy has three main advantages: 1) it enables our framework to adapt to various tasks; 2) it seamlessly supports CFG during inference; 3) it allows flexible management of the training process to avoid retraining due to errors. Meanwhile, this training strategy, from single modality to multiple modalities, can be considered as first learning the representation of each modality separately, and then learning the representation of the interaction between multiple modalities, which can be distinguished by the Task Instruction Modulation. Furthermore, by prioritizing the introduction of spatial conditions (i.e., source motion and trajectory, this strategy can share the model’s understanding between them and abstract conditions (i.e., text and style), as the latter conditions can be represented by the former to a certain extent.
@@ 5.4. MotionLab Inference
Ours 0.167 0.810 9.593 2.830 2.912 0.068
Table 3. Evaluation of text-based motion generation on HumanML3D(Guo et al., 2022a) dataset. The models in bold are the optimal models, and the models in underline are the sub-optimal models.
@@
Table 4. Evaluation of trajectory-based motion generation on HumanML3D (Guo et al., 2022a) dataset.
@@
Table 5. Evaluation of text-based and trajectory-based motion editing on MotionFix (Athanasiou et al., 2024) dataset. TMED mean that we re-implement the models since original models are trained on the skeleton of SMPL format, while our models are trained on HumanML3D format.
@@ 6. Experiments
We evaluate our framework using the following datasets. To evaluate the text-based motion generation, the trajectory-based motion generation, motion in-between and motion style transfer, we leverage the HumanML3D (Guo et al., 2022a) dataset, which comprises 14,646 motions and 44,970 motion annotations. To evaluate on the text-based and trajectory-based motion editing, we utilize MotionFix (Athanasiou et al., 2024) dataset, which is the first dataset for text-based human motion editing including 6,730 motion pairs.
@@
Evaluation Metrics. We evaluate our framework using the following metrics: 1) To evaluate text-based motion generation, following the (Chen et al., 2023), we introduce the FID to evaluate the distribution gap between the generated and original motions; Diversity to calculate the corresponding variance between motions; R-precision (R@K) to measure the proximity of the generated motion to the text or motion; Foot skating ratio to evaluate the physical plausibility of motion; Multi-modal Distance (MM Dist) calculates the distance between motions and texts. We also introduce Average Inference Time per Sample (AITS) measured in seconds to evaluate the inference efficiency; 2) To evaluate trajectory-based motion generation and motion in-between, following (Xie et al., 2023), we introduce the Average Error to measures the mean distance between the generated motion locations and the keyframe locations; 3) To evaluate text-based and trajectory-based motion editing, following (Athanasiou et al., 2024), we introduce the AvgR to measure the success rate of retrieval from edited motion to target motion; 4) To evaluate motion style transfer, following the (Song et al., 2024), we introduce the Style Recognition Accuracy (SRA) and Content Recognition Accuracy (CRA) to measure the stylistic and content accuracy of the generated motion; Trajectory Similarity Index (TSI) to evaluate the trajectory preservation from source motion.
Method text gen. (FID) traj. gen. (avg. err.) text edit (R@1) traj. edit (R@1) in-between (avg. err.) style transfer (CRA) style transfer (SRA)
w/o rectified flows 0.301 0.0359 54.38 69.21 0.0289 42.20 63.96
w/o MotionFlow Transformer 0.483 0.0447 51.26 65.34 0.0349 35.36 53.83
w/o Aligned ROPE 0.253 0.0886 45.39 61.99 0.0756 42.23 56.59
w/o task instruction modulation 0.223 0.0401 55.96 70.01 0.0288 40.55 63.91
w/o motion curriculum learning 1.956 0.1983 28.56 36.61 0.1682 29.51 34.23
Ours specialist models 0.209 0.0398 41.44 59.86 0.0371 43.53 67.55
Ours 0.167 0.0334 56.34 72.65 0.0283 44.62 69.21
Table 6. Ablation studies of key components of MotionLab on each task. Refer to the text for the detailed configuration of each variant.
@@
Implementation Details. In order to fairly compare our model with other models, motions from all datasets have been retargeted into one skeleton following HumanML3D format with 20 fps, where the number of joint $J$ is 22 and the dimension of motion feature $D$ is 263. The learning rate is set to be 1$\times$10-4. The timesteps are set to 1,000 for training and 50 for inference. Our models are trained by four RTX 4090D with each batch of 64 for 4 days. To ensure a fair comparison, the AITS of all models are recalculated using one RTX 4090D.
@@ 6.1. Quantitative Results
Overall Performance. As shown in Table 3 to Table 5, MotionLab demonstrates promising performance across all benchmarks [^1] Due to space limitations, we include the quantitative results on motion in-between and motion style transfer in the supplementary material., underscoring the effectiveness of our framework’s design. Notably, as MotionLab is a unified framework without task-specific designs, it must balance versatility, performance, and efficiency.
Specifically, as shown in Table 3 and Figure 7, MotionLab achieves superior performance (lowest FID, which is the key metric for generation tasks) with relatively fast inference time (third-lowest AITS). For trajectory-based tasks (Table 4) and the motion in-between task, MotionLab achieves lower average error. We believe these improvements stem from the effectiveness of masked pre-training and Aligned ROPE, which ensures spatial and temporal synchronization between the trajectory and target motion.
@@ 6.3. Ablation Studies
We perform several ablation experiments [^2] Additional ablation studies are available in the supp. material. on our framework to validate the designs in MotionLab and report the results in Table 6: the $1^{st}$ variant replaces rectified flows with diffusion models; the $2^{nd}$ variant uses a regular transformer (i.e., without modulation mechanism and adopting cross-attention) instead of MFT. The $3^{rd}$ variant uses the implicit 1D-learnable encoding instead of Aligned ROPE; The $4^{th}$ variant does not adopt the Task Instruction Modulation; the $5^{th}$ variant directly learns all tasks based on their FID compared to the last evaluation. Additionally, we use the same model to train specialist models for each task, denoted as ‘our specialist models’ in Table 6.
As can be seen from the results, the removal of motion curriculum learning markedly diminishes model performance across all tasks, underscoring its pivotal role in facilitating knowledge transfer between tasks. Meanwhile, our unified framework outperforms our specialist models in all tasks, potentially due to the knowledge sharing of motion curriculum learning. These phenomenons can be also attributed to the strategy’s capacity to enable the model to integrate its comprehension of spatial conditions (e.g., source motion, trajectory, and intermediate states) with abstract conditions (e.g., text and style), given that the latter can be partially represented by the former. Furthermore, as shown in Table 6 , Aligned ROPE is essential for space-related tasks, significantly reducing the average error. It effectively aligns source motion and target motion temporally, contributing to high R-precision in editing tasks.
Figure 7. Impact of timesteps during inference on MotionLab. The closer the model’s performance is to the lower left corner, the stronger the model is.
Additionally, we evaluate the impact of timesteps during inference on our MotionLab and compare its performance with baseline methods in terms of generation quality and inference time for the text-based motion generation task. As shown in Figure 7, our framework strikes an optimal balance between generation quality and efficiency.
@@ Appendix A Additional Quantitative Results
As shown in Table 7, our framework outperform CondMDI on all setting, illustrating the effectiveness of our framework in motion in-between.
[…]
CondMDI (Cohan et al., 2024) 1 0.1551 0.6787 9.5807 0.0936 0.3739
5 0.1731 0.6823 9.3053 0.0850 0.1789
20 0.2253 0.6821 9.1151 0.0806 0.0754
Ours 1 0.7547 0.6681 8.9058 0.0779 0.0875
5 0.0724 0.9146 9.4406 0.0504 0.0283
20 0.0288 0.9914 9.5447 0.0216 0.0215
Table 7. Evaluation of motion in-between with CondMDI (Cohan et al., 2024) on HumanML3D (Guo et al., 2022a) dataset.
Also as shown in Figure 8, our framework also outperform MCM-LDM on all metrics, demonstrating the effectiveness of our framework in motion style transfer.
@@ Appendix B Additional Ablation Studies
To further validate the Aligned ROPE, we also introduce the variant of 3D-Learnable and 3D-ROPE to distinguish the source motion, target motion and trajectory. As shown in Table 8 and Figure 9, 1D-position encoding are better than 3D-position encoding by avoiding introducing distances between different modalities, and ROPE are better than learnable position encoding by explicit positional encoding. Hence our 1D-ROPE outperforms all others variants, demonstrating its effective to embed the position information into tokens.
To further validate the motion curriculum learning, we adopt the variant of removing the masked pre-training and directly supervised fine-tuning in order; the variant of with masked pre-training but supervised fine-tuning all tasks together; the variant of introducing masked reconstruction, motion in-between and trajectory based motion generation in orderly. As shown in Table 9, our proposed motion curriculum learning outperforms all other variants, highlighting the effective of masked pre-training and fine-tuning tasks in order by avoiding gradient conflicts between different tasks. Specifically, the variant of masked pre-training in order demonstrates that necessity of introduce motion in-between and trajectory-based motion generation together, or will greatly weakens the performance of the model in the latter task.
To further validate the choice and combinations of the tasks, we also introduce the variants of different tasks. As shown in Table 10, improper combination of tasks will cause the unified framework to be weaker than the ours specialist models, while our carefully selected combination of all tasks makes our unified framework beat ours specialist models.
Figure 9. Ablation results of MotionLab on the motion in-between (with text). Beige motion is use 1D-learnable position encoding, purple motion use Aligned ROPE, and gray motions are the poses provided in keyframes, demonstrating the importance of Aligned ROPE.
Method text gen. (FID) traj. gen. (avg. err.) text edit (R@1) traj. edit (R@1) in-between (avg. err.) style transfer (CRA) style transfer (SRA)
1D-Learnable 0.246 0.0886 45.39 61.99 0.0756 39.40 56.59
3D-Learnable 0.346 0.1865 35.46 53.74 0.1460 36.99 58.81
3D-ROPE 0.241 0.0579 51.34 70.00 0.0354 42.96 62.46
1D-ROPE (ours) 0.167 0.0334 56.34 72.65 0.0273 44.62 69.21
@@
Method text gen. (FID) traj. gen. (avg. err.) text edit (R@1) traj. edit (R@1) in-between (avg. err.) style transfer (CRA) style transfer (SRA)
random selection based on FID 2.236 0.1983 28.56 36.61 0.1682 26.61 34.23
removing the masked pre-training 0.861 0.0932 44.99 63.92 0.0639 39.63 57.59
supervised fine-tuning all tasks together 1.331 0.1317 38.19 55.22 0.1143 36.60 50.59
masked pre-training in order 0.256 0.0423 56.33 69.31 0.0264 42.67 64.39
motion curriculum learning (ours) 0.167 0.0334 56.34 72.65 0.0273 44.62 69.21
@@
Task Metric
text gen. traj. gen text edit traj. edit in-between style transfer text gen. (FID) traj. gen. (avg. err.) text edit (R@1) traj. edit (R@1) in-between (avg. err.) style transfer (CRA) style transfer (SRA)
ours specialist models 0.209 0.0398 41.44 59.86 0.0371 43.53 67.55
$\times$ $\times$ $\times$ $\times$ 0.240 - - - - 41.23 65.53
$\times$ $\times$ $\times$ $\times$ 0.235 - 52.79 - - - -
$\times$ $\times$ $\times$ 0.176 0.0364 - - 0.0297 - -
$\times$ 0.171 0.0344 55.10 72.20 0.0287 - -
0.167 0.0334 56.34 72.65 0.0273 44.62 69.21
Table 10. Ablation studies of our MotionLab’s task combinations.